# Using $k$-medoids for distributed approximate similarity search with arbitrary distances

### v1

@@ 6 Conclusions and future work
This paper presents the GMASK algorithm for approximate similarity search in distributed systems that can accept arbitrary distance functions. One of the main novelties is adopting the $k$-medoids (PAM) algorithm, which is not bounded to the Euclidean distance, to achieve efficient solutions. Furthermore, using FastPAM, a highly efficient implementation of the PAM algorithm to solve the $k$-medoids problem, GMASK can improve its efficiency with respect to other previous algorithms.
Experiments conducted on real-world datasets with different properties confirm that GMASK, based on $k$-medoids, performs better in solving nearest-neighbour queries than other frequently used algorithms, using other Minkowski dissimilarity functions different from the Euclidean distance. These results illustrate that indexing data with non-Euclidean distances is not only viable but also holds the potential to enhance search outcomes substantially. Overlooking alternative distances would represent a voluntary limitation that may hinder us from achieving the optimal results.
GMASK lays out new interesting routes for further research. Once the requirement of using the Euclidean distance has been lifted, additional studies should be conducted to evaluate GMASK performance in other settings where alternative distances can be applied, such as pseudo-metrics (Minkowski distances with $p<1$) or non-metric dissimilarities. Besides, using arbitrary distances will enable us to consider datasets with qualitative features exclusively or a combination of quantitative and qualitative features. Many state-of-the-art algorithms and standard dissimilarity measures cannot be directly applied in these cases since they are incompatible with categorical features.

### v2

@@ 6 Conclusions and future work
This paper presents the GDASC algorithm for approximate similarity search in distributed systems that can accept arbitrary distance functions. One of the main novelties is the integration of $k$-medoids as clustering method, which is not bounded to the Euclidean distance, to achieve efficient solutions.
Experiments conducted on real-world datasets with different properties confirm that GDASC performs better in solving nearest-neighbour queries than other frequently used algorithms, using dissimilarity functions different from the Euclidean distance. These results illustrate that indexing data with non-Euclidean distances is not only viable but also holds the potential to enhance search outcomes substantially. Overlooking alternative distances would represent a voluntary limitation that may hinder us from achieving the optimal results.
GDASC lays out new interesting routes for further research. Once the requirement of using the Euclidean distance has been lifted, additional studies should be conducted to evaluate GDASC performance in other settings where alternative distances can be applied, such as pseudo-metrics (Minkowski distances with $p<1$) or non-metric dissimilarities. Besides, using arbitrary distances will enable us to consider datasets with qualitative features exclusively or a combination of quantitative and qualitative features. Many state-of-the-art algorithms and standard dissimilarity measures cannot be directly applied in these cases since they are incompatible with categorical features.
7 Availability of data and material
7.1 Data availability
All datasets used in experiments are publicly available:
7.2 Code availability
All code related to the proposed algorithm and the implementation of experiments is hosted in the elenagarciamorato/GDASC GitHub repository, specifically in the release with DOI 10.5281/zenodo.13220360).